# thanks: Jointly supervised this workthanks: Jointly supervised this work

### v1

@@ Abstract
Here, $\mathcal{\zeta}(Z,\theta)=(1+Z^{2})\cos^{2}{(\theta)}/[1+Z^{2}-\sin^{2}{(\theta)}]$ and $\gamma=\Gamma_{0}/\Delta\ll 1$ is a ‘small’ dimensionless parameter characterizing the tunneling strength. We seek solutions of Eq. (4) in powers of $\gamma$ [12]: $\left|\epsilon/\Delta\right|=c_{1}\gamma+c_{2}\gamma^{2}+c_{3}\gamma^{3}$, where $c_{1},c_{2}$ and $c_{3}$ are unknown complex coefficients. Then in the ‘short’ junction limit, we obtain two branches (labeled by $\pm$) of complex ABS spectrum with energies (see Appendix B for details):
$\frac{E^{(\pm)}_{b}(\phi)}{\Delta}$ $=$ ±$\pm\Bigg{[}(\gamma-\gamma^{2}+\gamma^{3})\left\{1-\mathcal{\zeta}(Z,\theta)\sin^{2}{\left(\frac{\phi}{2}\right)}\right\}^{\frac{1}{2}}$ (5)
$+$ $\frac{\gamma^{3}}{2}\left\{1-\mathcal{\zeta}(Z,\theta)\sin^{2}{\left(\frac{\phi}{2}\right)}\right\}^{\frac{3}{2}}\Bigg{]}\;.$
The real parts of $E^{(\pm)}_{b}$ represent physical energies of the ABSs while the imaginary part of $E^{(-)}_{b}$ represents their finite lifetime [31, 20]. The variation of the ABS spectrum with changing $\theta$ is shown in Fig. 2(a). For normal incidence ($\theta=0$), since $\zeta$ becomes independent of $Z$, the ABS spectra is gapless exhibiting a zero-energy crossing at $\phi=\pi$ and lack any imaginary components. This indicates no dissipation for normally incident quasiparticles preserving topological protection of the Klein tunneling effect [32] in 3DTI surface states [33, 34, 35]. In contrast, obliquely incident quasiparticles experience the effect of non-Hermiticity, resulting in two qualitatively new effects: finite lifetime for ABSs and a fully gapped ABS spectrum. The energy gap, $E_{\rm gap}(\phi)=\text{Re}[E^{(+)}_{b}(\phi)-E^{(-)}_{b}(\phi)]$, varies with $\theta$ and $Z_{i}$, as shown in Fig. 2(b).
The angle-resolved Josephson current $I(\theta,\phi)$ at zero temperature is calculated using the occupied part of the ABS spectrum with the formula [36, 37]:
$I(\theta,\phi)=\frac{2e}{\hbar}\frac{dE^{(-)}_{b}}{d\phi}\;,$ (6)
which for the NH-TIJJ gives,
$\frac{I(\theta,\phi)}{I_{0}}=\frac{\zeta\Big{[}1-\gamma+\frac{5}{2}\gamma^{2}-\frac{3}{2}\gamma^{2}\zeta\sin^{2}{\left(\frac{\phi}{2}\right)}\Big{]}}{\Big{[}1-\mathcal{\zeta}\sin^{2}{\left(\frac{\phi}{2}\right)}\Big{]}^{1/2}}\sin{(\phi)}\;.$ (7)
Here, $I_{0}=eN\Gamma_{0}/(2\hbar)$ with $N$ being the number of transmission channels. The variation of $I(\theta,\phi)$ for different $\theta$ is shown in Fig. 2(c). For normal incidence, $I(\theta,\phi)$ is always real as expected, $2\pi$-periodic and changes sign abruptly at $\phi=\pi$ due to the change in the ground state occupation at the zero-energy crossing (see Fig. 2(a)) [38, 39, 40]. On the other hand, for oblique incidence ($\theta\neq 0$), $I(\theta,\phi)$ becomes non-sinusoidal and complex. The magnitude of the real component decreases with increasing $\theta$ while the imaginary component increases, indicating that the effect of non-Hermiticity is more effective with increasing $\theta$.
The total equilibrium Josephson current ($I_{\rm tot}$) is obtained as:
[…]
Then the CPR for the NH-TIJJ is shown in Fig. 2(d). We note that the $2\pi$-periodic CPR for the Hermitian case ($Z_{2}=0$) that changes sign discontinuously at $\phi=\pi$ transforms into a non-sinusoidal CPR that changes sign smoothly at $\phi=\pi$ for the non-Hermitian case ($Z_{2}=0.1$).
Purely dissipative barrier: In order to demonstrate the sole effect of the imaginary potential barrier, we consider the $Z_{1}=0$ case which is the non-Hermitian counterpart of the Kulik-Omel’yanchuk limit with fully transparent barrier [41]. The corresponding ABS energy spectrum is shown in Fig. 3(a) for different $\theta$. To compare, we note that in the purely Hermitian case ($Z_{2}=0$, $Z_{1}\neq 0$), the ABS spectrum has a zero-energy crossing at $\phi=\pi$ at $\theta=0$ and it changes from gapless to gapped for $\theta\neq 0$ with gradually increasing gap for larger $\theta$ [12]. In contrast, for this purely non-Hermitian case, the ABS spectrum always remains gapless and the zero-energy crossing at $\phi=\pi$ for $\theta=0$ evolves into a line of ZES on the $\phi$-axis for any $\theta\neq 0$. The ZES also have a non-zero imaginary energy component indicating their finite lifetime. The boundary of the ZES-line is determined by the condition: $\mathcal{\zeta}(Z_{2},\theta)\sin^{2}{(\phi/2)}=1$ giving the position of the boundary as:
$\phi_{n}(Z_{2},\theta)=2n\pi\pm 2\sin^{-1}{(1/\sqrt{\zeta(Z_{2},\theta)})}\;,$ (9)
where, $n$ is an integer. These boundary points mimic the behavior of exceptional points usually observed in non-Hermitian systems [42, 22] and are tunable by $Z_{2}$ and $\theta$, as shown in Fig. 3(b). The blue region indicates the ZES and its boundary given by Eq. (9) characterizes the location of the exceptional-like points that widens as $Z_{2}$ increases, indicating that the strength of non-Hermiticity is directly responsible for the formation and tuning of the positions of these points (see also Appendix B for more details). However, for normal incidence we see that the spectrum (Fig. 3(a)) remains the same as in the Hermitian case and there is no formation of the exceptional-like points. This is expected since for normal incidence the effect of non-Hermiticity is not felt by the quasiparticles due to Klein tunneling, as discussed earlier.
To analyze the nature of the ZES further, we note that their wavefunctions for $\phi=(2n+1)\pi$, with $n$ being an integer, are (see Appendix B): $\Sigma_{\pm}(x,\theta)=\chi_{\pm}(\theta)e^{\pm ik_{F}x-x/\xi}$ with $\chi_{\pm}(\theta)=\left(1,\pm e^{\pm i\theta},-ie^{\pm i\theta},\pm i\right)^{T}$. These two wavefunctions are orthogonal to each other which prevents mixing of two zero energy ABSs. For normal incidence, they are also time reversal partners of each other as they are connected by the time reversal symmetry operation as $\Sigma_{-}(x,0)=\tau_{0}i\sigma_{y}\Sigma_{+}(x,0)$ [12] but this is not the case for a general $\theta\neq 0$. Hence, for $\theta=0$, the ZES are indeed Majorana zero modes protected by time-reversal symmetry [8] but for general $\theta\neq 0$, they are not Majorana modes.
The angle-resolved Josephson supercurrent in Eq. (7) diverges at the exceptional-like points as shown in Fig. 3(c) similar to other NH-JJs where non-Hermicity is introduced within the superconductor Hamiltonian [19] or through the self-energy term [18, 22]. Hence, we compute the total physical Josepshon current in Eq. (8) by excluding the exceptional-like points avoiding the divergences. The corresponding CPR is shown in Fig. 3(d). We note that the $2\pi$-periodic CPR for the fully transparent junction (green) becomes non-sinusoidal for this purely non-Hermitian (pink) case and the magnitude of the critical current also decreases.
@@
Here, $Z^{2}=(Z^{2}_{1}+Z^{2}_{2})$ with the dimensionless barrier parameters defined as $Z_{i}=\frac{mV_{i}}{\hbar^{2}k_{F}}$ in this case, $g_{1}(\theta,\phi)=Z^{2}-\cos^{2}{(\theta)}\sin^{2}{\left(\frac{\phi}{2}\right)}$, $g_{2}(\theta,\phi)={\left[Z^{2}+\cos^{2}{(\theta)}\right]^{2}-4Z^{2}_{2}\cos^{2}{(\theta})}$ and $\Delta^{\prime}$ is the induced superconducting gap amplitude in the metal. We note that the ABS solutions exist only in the allowed regime: $\phi\in[\phi_{l},\phi_{u}]$ in contrast to the NH-TIJJ case. Furthermore, the ABSs decay exponentially with a decay length $\lambda=\frac{\hbar v_{F}}{\Delta^{\prime}}(1-E^{2}_{b}/\Delta^{2})^{-1/2}$, which sets the condition for their existence: $\text{Re}[\lambda]>0$. This condition results in Josephson gaps, defined as intervals in $\phi$ where there are no ABS solutions. The phase boundaries of the Josephson gaps [20] are given by: $\phi_{l}=2n\pi+\phi_{0}(Z)$ and $\phi_{u}=(2n+1)\pi-\phi_{0}(Z)$ where $\phi_{0}(Z)=2\sin^{-1}\left({\frac{\sqrt{2}Z_{2}}{\sqrt{\cos^{2}{(\theta)}+Z^{2}}}}\right)$ with $n$ being an integer.
The ABS spectra for $\theta=0$ and a finite $\theta=\pi/6$ are shown in Fig. 4(a). The spectra is complex, with both the energy gap between the real branches and the behavior of the imaginary part strongly dependent on the angle of incidence. ${\rm Im}[E^{(-)}_{b}]$ increases with increasing $\theta$ which indicates that the effect of non-Hermiticity increases with increasing $\theta$ consistent with the NH-TIJJ case (Fig. 2). The boundary of the Josephson gaps (shaded regions in Fig. 4) also increases with increasing $\theta$ which demonstrates non-Hermiticity effect is controlled both by the imaginary potential strength $Z_{2}$ and $\theta$. Hence, the key differences between the ABS spectra of this ‘ordinary’ NH-JJ and the previously described NH-TIJJ are: (1) the presence of Josephson gaps in the NH-JJ, which are absent in the NH-TIJJ, and (2) the absence of zero-energy states in the NH-JJ, which were present in the NH-TIJJ.
Figure 4: Characteristics of an ‘ordinary’ NH-JJ: (a) The ABS spectrum of an ‘ordinary’ NH-JJ (the 3DTI in Fig. 1 is replaced by a metal) for different $\theta$. (b) Variation of the complex angle-resolved supercurrent for $\theta=0$, and the CPR of the junction. Here, $I^{\prime}_{0}=Ne\Delta^{\prime}/\hbar$ and $Z_{1}=Z_{2}=0.2$. The shaded regions indicate the Josephson gaps.
The variation of the real and imaginary parts of the the angle-resolved Josephson current $I(\theta,\phi)$ in Eq. (6) for normal incidence is shown in Fig. 4(b). We note that the absence of any Josephson current within the Josephson gaps as expected since there are no ABS solutions within these gaps. The total Josephson current ($I_{\rm tot}$) also shown in Fig. 4(d) demonstrates that the CPR is non-sinusoidal and exhibits a Josephson gap same as that of the $\theta=0$ case in Fig. 4(a). This is in stark contrast with the CPR for the NH-TIJJ case (shown in Fig. 3(d)), which exhibits no Josephson gaps.
Summary and outlook: We have investigated the quantum transport properties of a dissipative Josephson junction on the surface of a 3DTI, aiming to elucidate the interplay between dissipation and topological protection. Dissipation, modeled via an effective non-Hermitian Hamiltonian within the Lindblad formalism, arises from coupling to a ‘lossy’ metallic lead [18, 24] at the junction to an electron reservoir. Our study reveals that the Andreev bound state (ABS) spectrum is generally complex for oblique incidence, indicating their finite lifetimes and decoherence in Cooper pair transport due to non-Hermiticity in this NH-TIJJ. Notably, normally incident quasiparticles preserve Klein tunneling, maintaining a real gapless ABS spectrum. Unlike ‘ordinary’ NH-JJs, NH-TIJJs lack Josephson gaps, highlighting their topological protection. Moreover, in the purely non-Hermitian regime, a tunable line of ZES emerges along the phase difference axis, with boundary points exhibiting exceptional-point-like features such as divergent angle-resolved supercurrents. Non-Hermiticity also reshapes the CPR, transforming it from a discontinuous $2\pi$-periodic form into a smoothly varying, non-sinusoidal function while preserving $2\pi$-periodicity.
@@
The physics of the ‘short’ NH-TIJJ described in this paper can be realized in practical devices, such as Nb-superconductors on the surface of a HgTe 3DTI [43], with dissipation introduced by attaching a ‘lossy’ metallic lead [18, 24] at the junction. For such a setup, relevant parameters include $\Delta\simeq 1$ meV, $\Gamma_{0}\simeq 0.2$ meV, $\hbar v_{F}\simeq 250-350$ meV-nm and the junction width $<1.25-1.75$ $\mu$[43]. Another promising platform is Nb-Bi2Te3-Nb JJs [44, 45]. Our findings thus highlight the intricate interplay between topology, superconductivity, and dissipation, providing critical insights into the experimental realization of dissipation-engineered hybrid quantum devices and paving the way for leveraging non-Hermitian effects in TIJJs to advance quantum computing architectures.

### v2

@@ Abstract
where $\gamma=\Gamma_{0}/\Delta\ll 1$ is a dimensionless parameter characterizing the tunneling strength. We seek solutions of Eq. (6) in powers of $\gamma$ [Tkachov2013]: $\left|\epsilon/\Delta\right|=c_{1}\gamma+c_{2}\gamma^{2}+c_{3}\gamma^{3}$, where $c_{1},c_{2}$ and $c_{3}$ are unknown complex coefficients. Then in the ‘short’ junction limit, we obtain two branches (labeled by $\pm$) of complex ABS spectrum with energies (see Appendix B for details):
$\frac{E^{\pm}_{b}(\theta,\phi)}{\Delta}=\pm\Bigg[(\gamma-\gamma^{2}+\gamma^{3})\mathcal{S}^{\frac{1}{2}}(\theta,\phi)+\frac{\gamma^{3}}{2}\mathcal{S}^{\frac{3}{2}}(\theta,\phi)\Bigg]\;.$ (7)
The real parts of $E^{\pm}_{b}$ represent physical energies of the ABSs while the negative imaginary part of $E_{b}^{\pm}$ represents the broadening of the energy levels due to non-Hermiticity [Capecelatro2025, Li2024], the positive imaginary branch signifies pumping of quasiparticles from the reservoir [Kornich2022] and therefore does not represent a physical solution for dissipative systems. The variation of the ABS spectrum for different angles of incidence $\theta$ is shown in Fig. 2(a). Before discussing the features introduced by non-Hermiticity, we first remind ourselves of the ABS characteristics in the Hermitian case; the spectrum for normal incidence ($\theta=0$) is gapless and features a zero-energy crossing at $\phi=\pi$ which corresponds to two Majorana bound states whose spinors are related by time-reversal symmetry and also preserve Klein tunneling [Tkachov2013]. The spectrum for oblique incidence ($\theta\neq 0$), however, features an angle-dependent gap and lacks time-reversal symmetry protection and thus experience scattering by the potential barrier.
In the non-Hermitian case, the normally incident quasiparticles feature a spectrum whose real part is gapless but the zero-energy crossing is shifted away from $\phi=\pi$ in contrast to the Hermitian case. There is also a finite imaginary component accompanying the real counterpart which indicates that Klein tunneling, which exists for normal incidence in the Hermitian case [Tkachov2013], is not robust against dissipation. The shift of the zero-energy crossing along the $\phi$ axis makes the spectrum asymmetric with respect to $\phi=\pi$ over the interval $\phi\in[0,2\pi]$. Moreover, the finite imaginary component introduces a finite lifetime for the bound states, rendering the associated Majorana bound states unstable in the presence of non-Hermiticity. This can be understood as a purely non-Hermitian effect where dissipation affects the Kramers degeneracy and can induce effects such as splitting of spectral functions of Kramers degenerate states [Deng2021].
Meanwhile, obliquely incident quasiparticles experience the effect of non-Hermiticity in two qualitatively new effects: (i) broadening of the quasi-ABSs energy levels which means that the bound states have acquired a finite lifetime [datta1997electronic] and (ii) a fully gapped quasi-ABS spectrum which beyond a certain interval $\phi\in[\phi_{1},\phi_{2}]$, the real part of both particle and hole spectrum bifurcates into two branches with one branch going above the induced gap (supra-gap) and the other going below it (sub-gap). We call the points $\phi_{1}$ and $\phi_{2}$ on the $\phi$-axis where the real part of the spectrum exits the gap as gap-exit points (GEPs), the imaginary part of the spectrum vanishes beyond these points as shown in Fig. 2(b) and the position of the GEPs, demonstrated by the boundary of the dark purple region, is asymmetric about $\phi=\pi$ in the interval $\phi\in[0,2\pi]$. Since the normally incident quasi-particles do not experience scattering from the real barrier, therefore the bifurcation is demonstrated only in oblique incidence and is absent for normal incidence.
Figure 3: (a) Variation of the imaginary part of the Andreev spectrum as a function of the real barrier parameter $Z_{1}$ and phase difference $\phi$. (b) The total current plotted as a function of $\phi$ for different values of the barrier parameter $Z_{1}$. The imaginary barrier parameter is fixed at $Z_{2}=0.2$ in both cases.
Figure 4: Characteristics of the ordinary NH-$J$ for a general complex barrier ($Z_{1}\neq 0,Z_{2}\neq 0$): (a) The real (solid lines) and imaginary (dashed lines) parts of the Andreev spectrum of the 2D NH-$J$ as a function of the phase difference $\phi$ plotted for different angles of incidence $\theta$ taking the imaginary barrier parameter $Z_{2}=0.2$. (b) Variation of the imaginary part of the spectrum as a function of $\phi$ and $Z_{2}$ for the incident angle $\theta=\pi/3$, the boundary of the dark purple region demonstrates the symmetric gap-exit points in the spectrum. (c) The angle-resolved current $I_{\theta}(\phi)$ plotted as a function of $\phi$ for different angles of incidence at zero temperature. (d) The total supercurrent $I_{\mathrm{total}}(\phi)$ as a function of $\phi$ for different barrier parameters. Other parameters are the same as given in Fig. 2.
To calculate the Josephson current, we employ the Furusaki-Tsukada (FT) formula [Furusaki1991, Lu2018] which provides an alternative approach to compute the supercurrent in terms of the Andreev reflection processes as,
$I_{\theta}(\phi)=\frac{e\Delta}{\hbar}k_{B}T\sum_{\omega_{n}}\frac{1}{\Omega_{n}}[r_{he}(\phi,\omega_{n})-r_{eh}(\phi,\omega_{n})]\;,$ (8)
where analytic continuation for the energy is done as $E\to i\omega_{n}$ in terms of the Matsubara frequencies $\omega_{n}=(2n+1)\pi/k_{B}T$ with $n$ being an integer, the parameter $\Omega_{n}=\sqrt{\omega^{2}_{n}+\Delta^{2}}$, $r_{he}$ is the Andreev coefficient for an incoming electron-like quasiparticle being reflected as a hole-like quasiparticle and $r_{eh}$ is the corresponding coefficient for the reverse process for a hole-like quasiparticle. In the zero-temperature limit $T\to 0$, the summation over Matsubara frequencies can be replaced by an integration over real energy as $k_{B}T\sum_{\omega_{n}}\to\frac{1}{2\pi}\int d\omega$ on performing the analytic continuation $i\omega_{n}\to\omega+i0^{+}$. We note that the existing supercurrent formulae for NH systems used to evaluate the current in systems with exceptional points [Pino2025, Shen2024] incur a divergence of the angle-resolved current at the GEPs, therefore to resolve these divergences we use the FT formula and limit our analysis within the subgap approximation ($|\omega|<\Delta$). The variation of the angle-resolved current $I_{\theta}(\phi)$ for different $\theta$ at zero temperature within the subgap approximation is shown in Fig. 2(c). The current is $2\pi$-periodic and exhibits an ‘anomalous’ behavior where the supercurrent is non-zero at $\phi=0,2\pi$ for all angles of incidence. For normal incidence, the current changes sign smoothly and does not exhibit the abrupt discontinuity at the zero-energy crossing as in the Hermitian case, which highlights the contribution of the imaginary part of the spectrum in reshaping and smoothing the CPRs. For oblique incidence, in addition to the mentioned characteristics, the critical current decreases with increasing $\theta$ indicating that the effect of non-Hermiticity is more effective with increasing obliqueness.
The total equilibrium Josephson current ($I_{\rm total}$) can be obtained as:
$I_{\text{total}}(\phi)=\int_{-\pi/2}^{\pi/2}I(\phi,\theta)\cos{(\theta)}d\theta\;.$ (9)
The CPR for the NH-TIJJ is shown in Fig. 2(d) for different values of the non-Hermiticity strength $Z_{2}$. The CPR features a $2\pi$-periodic, non-sinusoidal, anomalous current which is asymmetric about $\phi=\pi$ carrying the signature of the asymmetry of the ABSs shown in Fig. 2. The critical current is maximum for the Hermitian case ($Z_{2}=0$) and decreases with increasing $Z_{2}$ for a fixed $Z_{1}$ due to quasiparticles experiencing decoherence and escaping to the reservoir. The main feature of the CPR is its asymmetry about the phase difference $\phi=\pi$ and its non-zero value at $\phi=0,2\pi$, this anomalous behavior of the supercurrent stems from the asymmetry in the Andreev spectrum whereby the corresponding subgap contribution outside the GEPs is also asymmetric, thus resulting in an anomalous CPR. The asymmetry is an outcome of the interplay between the real barrier and non-Hermiticity, as is evident in Fig. 3(a) where the GEPs (boundary of dark purple region) are initially symmetric when $Z_{1}=0$ and become asymmetric when $Z_{1}\neq 0$. The variation of the CPR for different values of the real barrier $Z_{1}$ for a fixed $Z_{2}$ is shown in Fig. 3(b), the critical current is maximum for $Z_{1}=0$ and decreases as $Z_{1}$ increases due to enhanced normal reflection. Note that the anomalous behavior is absent when $Z_{1}=0$ since asymmetry in the spectrum requires both the barrier parameters $Z_{1}$ and $Z_{2}$ to be non-zero.
@@
Here, $Z^{2}=(Z^{2}_{1}+Z^{2}_{2})$ with the dimensionless barrier parameters defined as $Z_{i}=\frac{mV_{i}}{\hbar^{2}k_{F}}$ in this case, $g_{1}(\theta,\phi)=Z^{2}-\cos^{2}{(\theta)}\sin^{2}{\left(\frac{\phi}{2}\right)}$, $g_{2}(\theta,\phi)={\left[Z^{2}+\cos^{2}{(\theta)}\right]^{2}-4Z^{2}_{2}\cos^{2}{(\theta})}$ and $\Delta^{\prime}$ is the induced superconducting gap amplitude in the metal.
The full Andreev spectrum is shown in Fig. 4 and it shares features similar to that of the NH-TIJJ. The spectrum for different incident angles $\theta$ is shown in Fig. 4(a). The spectrum is complex with the real branches showing bifurcation beyond the interval $\phi\in[\phi_{0},2\pi-\phi_{0}]$ where $\phi_{0}(Z,\theta)=2\sin^{-1}\left({\frac{\sqrt{2}Z_{2}}{\sqrt{\cos^{2}{(\theta)}+Z^{2}}}}\right)$, both the energy gap between the real branches and the behavior of the imaginary part are strongly dependent on the angle of incidence. ${\rm Im}[E_{b}]$ increases with increasing $\theta$ which indicates that the effect of non-Hermiticity increases with increasing $\theta$ consistent with the NH-TIJJ case (see Fig. 2). The boundary of the purple region (see Fig. 4(b)) where the imaginary part of the spectrum for $\theta=\pi/3$ is plotted as a function of the imaginary potential strength $Z_{2}$ and $\theta$ demonstrates the position of the GEPs which are symmetric about $\phi=\pi$ in contrast to the asymmetric feature in NH-TIJJ (see Fig. 2(b)). Hence, the NH-JJ and NH-TIJJ share a common qualitative behavior in that the Andreev spectrum is complex with the real parts exhibiting bifurcation beyond the GEPs while also exhibiting contrasting behaviors such as the absence of zero-energy states in an NH-JJ and an asymmetric spectrum in the NH-TIJJ.
The variation of the angle-resolved Josephson current $I_{\theta}(\phi)$ in Eq. (8) for different angles of incidence is shown in Fig. 4(c). We note that in stark contrast to the NH-TIJJ, the angle-resolved current does not exhibit the anomalous behavior since the Andreev spectrum of the NH-JJ is symmetric. The total current shown in Fig. 4(d) for different values of $Z_{2}$ features a $2\pi$-periodic, non-sinusoidal CPR whose critical current decreases with increasing $Z_{2}$. The absence of anomalous behavior of the supercurrent in the case of NH-JJ is due to the lack of any Kramers degeneracy in the spectrum, thus there is no interplay with non-Hermiticity to create the asymmetry in the spectrum as in the case of a NH-TIJJ.
Figure 5: Comparison of the CPR computed from the Furusaki-Tsukada (FT) formula (Eq. (8)) and the Andreev-bound-state (ABS) formula (Eq. (11)) for: (a) planar ordinary NH-$J$. (b) strong-proximity NH-TIJJ. The barrier parameters are fixed at $Z_{1}=0.4,Z_{2}=0.2$ in both cases.
Comparison of CPR from ABS formula: Given the inability of the alternative supercurrent formulae for NH systems with EPs to resolve the divergence of the supercurrent at the GEPs in our system as mentioned earlier, we calculate the supercurrent from the ABS formula derived from the polar Green’s function [Capecelatro2025] and compare with that computed from the FT formula. For this purpose, we numerically compute the angle-resolved Josephson current at zero temperature as [Capecelatro2025]:
$I_{\theta}(\phi)=$ $-\frac{2e}{\hbar}\frac{1}{\pi}\Big\{(\partial_{\phi}\epsilon)\int_{-\Delta}^{\Delta}d\omega\,\frac{\tanh(\beta\omega/2)\,\lambda}{(\omega+\epsilon)^{2}+\lambda^{2}}$ (11)
$-(\partial_{\phi}\lambda)\int_{-\Delta}^{\Delta}d\omega\,\frac{\tanh(\beta\omega/2)\,(\omega+\epsilon)}{(\omega+\epsilon)^{2}+\lambda^{2}}\Big\}$
after decomposing the Andreev spectrum as $E_{b}=\epsilon-i\lambda$ with $\epsilon$ and $\lambda$ being the real and imaginary part of the spectrum respectively. The resulting CPR is plotted in Fig. 5 for both the NH-TIJJ and NH-JJ. Within the subgap approximation, the CPR computed from the polar ABS formula exhibits intervals along $\phi$, where the supercurent vanishes, called Josephson gaps. The appearance of Josephson gaps outside the GEPs is very similar to the results shown in Ref. [Li2024]. However, notice that there is a sizable discrepancy between the two approaches, more prominently around the GEPs, there are no Josephson gaps in the CPR computed using the FT formula. This mismatch arises because the ABS formula takes into account only the contribution from the bound-state spectrum which comes from the poles of the Green’s function, thereby resulting in the absence of supercurrent outside the GEPs as there are no bound states. However, in non-Hermitian systems, there is another contribution stemming from the branch-cut part of the Green’s function [Capecelatro2025]. This latter contribution is not captured by the polar ABS formula but is included in the FT formula since the current is computed by taking into account all the reflection processes, thus giving a finite supercurrent outside the GEPs.
Summary and conclusions: We have investigated the quantum transport properties of a dissipative Josephson junction on the surface of a 3DTI, aiming to elucidate the interplay between dissipation and topological protection. Dissipation, modeled via an effective non-Hermitian Hamiltonian within the Lindblad formalism, arises from coupling to a ‘lossy’ metallic lead [cayao2023non, Ohnmacht2024] at the junction to an electron reservoir. Our study reveals that the Andreev bound state (ABS) spectrum is complex and merges into the continuum regime beyond certain interval of the phase difference. The spectrum also exhibit an asymmetry arising from the interplay of normal scattering and non-Hermiticity. This asymmetry in the spectrum results in the anomalous behavior of the CPR, where a finite supercurrent is present even when the phase difference is zero.
@@
The physics of the ‘short’ NH-TIJJ described in this paper can be realized in practical devices, such as Nb-superconductors on the surface of a HgTe 3DTI [Maier2012], with dissipation introduced by attaching a ‘lossy’ metallic lead [cayao2023non, Ohnmacht2024] at the junction. For such a setup, relevant parameters include $\Delta\simeq 1$ meV, $\Gamma_{0}\simeq 0.2$ meV, $\hbar v_{F}\simeq 250-350$ meV-nm and the junction width <1.251.75<1.25-1.75 $\mu$[Maier2012]. Another promising platform is Nb-Bi2Te3-Nb JJs [Stolyarov2022, Charpentier2017]. Since the length of the TIJJ qualitatively changes the CPR [Lu2023], a natural future direction is to investigate the effect of dissipation on quantum transport in a ‘long’ NH-TIJJ which will be explored in a separate study. Our findings thus highlight the effect of dissipation on the Josephson current in a topological Josephson junction, providing critical insights into the experimental realization of dissipation-engineered hybrid quantum devices and paving the way for leveraging non-Hermitian effects in TIJJs to advance quantum computing architectures.